OpenStudy (anonymous):

Simplify the radical. 7 square root of 72

OpenStudy (anonymous):

theses are the possible answers 42 square root of 2 63 square root of 8 21 square root of 2 21 square root of 8 @PinkSapphire

OpenStudy (jhannybean):

\[7\sqrt{72} = 7\sqrt{9\cdot 8} =7\sqrt{3^2 \cdot 2^2 \cdot 2}\] For every pair of 2 numbers that are alike, 1 comes out of the square root andmultiplies to the number outside of it. \[=\boxed{7 \cdot 3 \cdot 2\sqrt{2} = ~?}\]

OpenStudy (pinksapphire):

\[7\sqrt{72}\]Find the greatest factors of 72. \[7\sqrt{36\times2}\]36 is a perfect square so you cat take it out of the square root and it will become 6. \[7\times 6\sqrt{2}\]Now multiply 6 and 7.\[42\sqrt{2}\]